# The Assouad dimension of Kakeya sets in $\mathbb{R}^{3}$

### v1

@@ 4.1 Arrangements of rectangular prisms
i.e. $\bigcup_{R\in\mathcal{R}}Y(R)$ has discretized Assouad dimension $\geq 3-\varepsilon$ at scale $\delta$ and scale separation $\delta^{-\tau_{i}}\geq\delta^{-\eta}$. If this occurs, then we halt the iteration and conclude the proof.
Case (B). Next, suppose that Conclusion (B) holds. Let $s_{i+1}=s^{\prime}$; $t_{i+1}=t_{i}^{\prime}$; let $\mathcal{R}_{i+1}$ be the set of $s_{i+1}\times t_{i+1}\times 1$ prisms, with shading $Y_{i+1}=Y_{i}^{\prime}$, as described by Lemma 4.4. We have that $\mathcal{R}_{i+1}$ satisfies the Convex Wolff Axioms with error $\delta^{-\tau_{i+1}}$, and $Y_{i+1}$ is $\delta^{\tau_{i+1}}$ dense. Furthermore, the shading $Y_{i+1}$ satisfies (4.21) with $i+1$ in place of $i$.
Case (B.i). Suppose that Conclusion (B) holds and $t_{i+1}=1$. Recall that $s_{i+1}\leq\delta^{\varepsilon/2}.$ Thus we can apply Lemma 4.3 with $s_{i+1}$ in place of $\delta$. We have
@@
We can select $\tau_{i+1}$ and $\delta_{0}$ sufficiently small so that the LHS of (4.22) has size at least $s_{i+1}^{\varepsilon}|B(0,1)|$. Since the set on the LHS of (4.22) is contained in $B(0,1)$, we conclude that $\bigcup_{R\in\mathcal{R}}Y(R)$ has discretized Assouad dimension $\geq 3-\varepsilon$ at scale $\delta$ and scale separation $\delta^{-\varepsilon/2}\geq\delta^{-\eta}.$ If this occurs, then we halt the iteration and conclude the proof.
Case (B.ii). Suppose $t_{i+1}<1$; then $s_{i+1}/t_{i+1}=s_{i}/t_{i}$, so in particular $s_{i+1}\leq\delta^{\varepsilon}t_{i+1}$. Thus we have verified all of the properties required to continue the iteration, with $i+1$ in place of $i$. Note that $s_{i+1}\geq\delta^{-\varepsilon^{2}/2}s_{i}$; $t_{i+1}\geq\min(1,\delta^{\varepsilon}s_{i+1})$, and the iterative process halts if $t_{i+1}=1$. We conclude that this iterative process will halt after at most $N=\lfloor 2/\varepsilon^{2}\rfloor$ steps. ∎
@@ 4.2 Arrangements of tubes that cluster into rectangular prisms
In Section 4.1, we showed that unions of rectangular prisms that are far from being tubes (i.e. prisms of dimensions $s\times t\times 1$, with $s$ much smaller than $t$) have large discretized Assouad dimension. In this section, we will use these results to show that if an arrangement of $\delta$-tubes clusters into prisms that are far from being (fatter) tubes, then the union of these tubes must have large discretized Assouad dimension. This is made precise in Lemmas 4.8 and 4.10 below. We begin with a technical lemma.
@@
For all $\varepsilon>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $\rho\in[20\delta,1]$ and let $K\geq 1$. Let $(\mathbb{T},Y)_{\delta}$ be a set of essentially distinct $\delta$-tubes, with #$\#\mathbb{T}\geq K\rho/\delta$, and $Y$ a uniformly $\delta^{\eta}$-dense shading (see Definition 2.1). Let $R_{0}$ be a $10\delta\times\rho\times 1$ rectangle, and suppose each $T\in\mathbb{T}$ is contained in $R_{0}$.
@@
where the maximum is taken over all pairs of tubes $T,T^{\prime}$ intersecting $x$. We define $\operatorname{L-spread}(x)=0$ if at most one tube intersects $x$. Let $C$ be a large number to be determined below. After dyadic pigeonholing, we can select $\theta\in[\frac{K\delta}{C},\rho]$ so that the shading
@@
(Note that the shading $Z(T)$ is not necessarily a subset of $Y(T)$). Our argument that such a shading $Z(T)$ must exist proceeds as follows: if we could not select a $\theta\in[\frac{K\delta}{C},\rho]$ so that (4.25) holds, then the shading
@@
satisfies $\sum_{T\in\mathbb{T}}|Z^{\prime}(T)|\geq\frac{1}{100}\delta^{2}(\#\mathbb{T})$. But for each $x\in\mathbb{R}^{3}$, the set of tubes $T\in\mathbb{T}$ with $x\in Z(T)$ are contained in the $10\delta$-neighborhood of a plane (since they are contained in $R_{0}$), and also make angle $\leq\frac{K\delta}{C}$ with a common unit vector; since the tubes in $\mathbb{T}$ are essentially distinct, we conclude that $O(K/C)$ tubes $T\in\mathbb{T}$ can satisfy $x\in Z^{\prime}(T)$, and hence
@@
$\mathbb{T}_{1}=\{T\in\mathbb{T}\colon|Z(T)|\gtrsim|\log\delta|^{-1}|T|\}.$ (4.27)
Selecting the implicit constant in (4.27) appropriately, we can ensure that #$\#\mathbb{T}_{1}\gtrapprox\#\mathbb{T}$. For each $T_{1}\in\mathbb{T}_{1}$, let $x_{2},\ldots,x_{N},\ N\gtrapprox\theta/\delta$ be a set of $\delta/\theta$-separated points from $Y_{1}(T_{1})$. For each $x_{i}$, select $T_{i}\in\mathbb{T}$ with $x_{i}\in T_{i}$ and $\angle(T_{1},T_{i})\in[\theta,2\theta)$. A $L^{2}$ argument (see e.g. (4.11)) plus the fact that $|Y(T_{i})|\geq\delta^{\eta}$ shows that
@@
On the other hand, the set on the LHS of (4.28) is contained in $R_{0}\cap N_{2\theta}(T_{1})$, which is essentially a rectangular prism of dimensions $10\delta\times 4\theta\times 1$. Denote this prism by $R(T_{1})$. We have just shown that
@@
To conclude the proof, let $\omega=4\theta$ and let $\mathcal{R}_{1}=\{R(T_{1})\colon T_{1}\in\mathbb{T}_{1}\}$. We claim that (4.23) holds for $\mathcal{R}_{1}$. Indeed, this follows from the fact that $T_{1}\subset R(T_{1})$ for each $T_{1}\in\mathbb{T}_{1}$, and #$\#\mathbb{T}_{1}\gtrapprox\#\mathbb{T}$. By (4.29) we see that (4.24) holds for each $R\in\mathcal{R}_{1}$, provided we select $\eta$ and $\delta_{0}$ sufficiently small. To conclude the proof, greedily select a set $\mathcal{R}\subset\mathcal{R}_{1}$ of essentially distinct rectangular prisms; we can perform this selection so that (4.23) continues to hold. ∎
@@
For all $\varepsilon>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $\mathbb{T}$ be a set of essentially distinct $\delta$-tubes that satisfies the Convex Wolff Axioms with error $\delta^{-\eta}$. Let $Y$ be a shading with $|Y(T)|\geq\delta^{\eta}|T|$ for each tube $T$. Then at least one of the following must occur.
@@
    There is a set $\mathbb{T}^{\prime}\subset\mathbb{T}$ with #$\#\mathbb{T}^{\prime}\geq\frac{1}{2}(\#\mathbb{T})$ so that for every $\delta\leq\rho\leq 1$ and every $\delta\times\rho\times 1$ rectangle $R$, at most $\delta^{-\varepsilon}(\rho/\delta)$ tubes from $\mathbb{T}^{\prime}$ are contained in $2R$.
@@
    $\bigcup Y(T)$ has discretized Assouad dimension at least $3-\varepsilon$ at scale $\delta$ and scale separation $\delta^{-\eta}$.
@@
Since $\mathbb{T}$ satisfies the Convex Wolff Axioms, we have that at most $\delta^{-\eta}(\rho\delta)(\#\mathbb{T})$ tubes are contained in each $2\delta\times 2\rho\times 2$ rectangle. If #$\#\mathbb{T}$ is substantially larger than $\delta^{-2}$, then the estimate from Conclusion (A) is stronger.
@@
We say a $2\delta\times 2\rho\times 2$ rectangle is “heavy” if it contains more than $\delta^{-\varepsilon}(\rho/\delta)$ tubes from $\mathbb{T}$. If fewer than $\frac{1}{2}(\#\mathbb{T})$ of the tubes are contained in heavy rectangles, let $\mathbb{T}^{\prime}$ be the set of tubes that are not contained in heavy rectangles. Then Conclusion (A) holds, and we are done.
Suppose instead that at least $\frac{1}{2}(\#\mathbb{T})$ tubes are contained in heavy rectangles. Since the tubes in $\mathbb{T}$ are essentially distinct, each $2\delta\times 2\rho\times 2$ rectangle can contain at most $O\big{(}(\rho/\delta)^{2}\big{)}$ tubes from $\mathbb{T}$. Thus if a $2\delta\times 2\rho\times 2$ rectangle is heavy, then we must have $\rho\gtrsim\delta^{1-\varepsilon}$. Let $\mathcal{R}_{0}$ denote this set of heavy rectangles (note that the rectangles in $\mathcal{R}_{0}$ can have different dimensions). Let $\varepsilon_{1}$ be a small quantity to be chosen below. Apply Lemma 4.7 to each rectangle $R_{0}\in\mathcal{R}_{0}$, with $\varepsilon_{1}$ in place of $\varepsilon$, and $K=\delta^{-\varepsilon}$. We obtain a number $\omega(R_{0})\gtrsim\delta^{1-\varepsilon}$ and a set $\mathcal{R}(R_{0})$ of $\delta\times\omega(R_{0})\times O(1)$ rectangles. We have
@@
Since $\mathbb{T}$ satisfies the Convex Wolff Axioms with error $\delta^{-\eta}$, by (4.30) we have that the set (i.e. not multi-set) $\mathbb{T}_{2}$ satisfies the Convex Wolff Axioms with error $\lessapprox\delta^{-\eta}$. Finally, for each $T\in\mathbb{T}_{2}$, let $R(T)$ be a rectangle in $\mathcal{R}_{2}$ with $T\subset R$ (if more than one $R\in\mathcal{R}_{2}$ contains $T$, simply select one of them). Define the multi-set =$\mathcal{R}=\{R(T)\colon T\in\mathbb{T}_{2}\}$, i.e. $\mathcal{R}$ is a multi-set with the same cardinality as $\mathbb{T}_{2}$. By Remark 2.7, we conclude that $\mathcal{R}$ satisfies the Convex Wolff Axioms with error $\lessapprox\delta^{-\eta}$.
@@
By Lemma 4.7, we have $|Y(R)|\geq\delta^{\varepsilon_{1}}|R|$ for each $R\in\mathcal{R}$. To conclude the proof, apply Proposition 4.6 with $\varepsilon$ as above, $s=\delta$, and $t=\omega_{0}$; to ensure that Proposition 4.6 can be applied, we must select $\varepsilon_{1}$ sufficiently small depending on $\varepsilon$ (our choice of $\varepsilon_{1}$ depends on the quantity “$\eta$” from Proposition 4.6), and then $\eta$ sufficiently small depending on $\varepsilon_{1}$. Since $\bigcup Y(T)\subset\bigcup Y(R),$ we conclude that Conclusion (B) holds. ∎
Lemma 4.8 says that either a union of $\delta$-tubes has large discretized Assouad dimension, or these tubes cannot concentrate into $\delta\times\rho\times 1$ rectangular prisms. The next lemma is similar, except the latter conclusion is strengthened: either a union of $\delta$-tubes has large discretized Assouad dimension, or these tubes cannot concentrate into $s\times t\times 1$ rectangular prisms, for any $\delta\leq s\leq t\leq 1$. To prove this result, we apply Lemma 4.8 to the $s$-thickening of the tubes in our arrangement, for many different values of $s$. The precise statement is as follows.
@@
For all $\varepsilon=1/N>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $(\mathbb{T},Y)_{\delta}$ be a set of essentially distinct $\delta$-tubes that satisfies the Convex Wolff Axioms with error $\delta^{-\eta}$, and let $Y$ be a uniformly $\delta^{\eta}$-dense shading. Then at least one of the following must occur.
@@
    There is a set $\mathbb{T}^{\prime}\subset\mathbb{T}$ with #$\#\mathbb{T}^{\prime}\geq\delta^{\varepsilon}(\#\mathbb{T})$ so that the following holds. For every pair of numbers $\delta\leq s\leq t\leq 1$ with $s$ of the form $\delta^{i\varepsilon}$, there is a 1-uniform partitioning cover $\mathbb{T}_{s}^{\prime}$ of $\mathbb{T}^{\prime}$ (see Definition 2.10), where $\mathbb{T}_{s}^{\prime}$ is a set of essentially distinct $s$-tubes. Furthermore, for every $s\times t\times 1$ rectangular prism $R$, at most $\delta^{-\varepsilon}\frac{t}{s}\frac{(\#\mathbb{T}^{\prime})}{(\#\mathbb{T}^{\prime}_{s})}$ tubes from $\mathbb{T}^{\prime}$ are contained in $R$.
@@
    $\bigcup Y(T)$ has discretized Assouad dimension at least $3-\varepsilon$ at scale $\delta$ and scale separation $\delta^{-\eta}$.
@@
Our first step is to regularize the set $\mathbb{T}$ at scales of the form $s_{i}=\delta^{i/N}$, with $i=1,\ldots,N$ (we do not need to consider $s_{0}=1$, since in that case there is nothing to prove). After repeated pigeonholing, we can select a set $\mathbb{T}^{\dagger}\subset\mathbb{T}$ with #$\#\mathbb{T}^{\dagger}\gtrsim(C|\log\delta|)^{N}(\#\mathbb{T})$ and sets of $s_{i}$-tubes $\mathbb{T}_{s_{i}}^{\dagger},\ i=0,\ldots,N$, so that $\mathbb{T}_{s_{N}}^{\dagger}=\mathbb{T}^{\dagger}$, and each set $\mathbb{T}_{s_{i}}^{\dagger}$ is a partitioning cover of $\mathbb{T}_{s_{i+1}}^{\dagger}$ (and hence also a partitioning cover of $\mathbb{T}_{s_{j}}^{\dagger}$ for all $j>i$). See e.g. Lemma 3.4 from [11] for an explicit description of this iterated pigeonholing process. We will choose $\delta_{0}$ sufficiently small so that #$\#\mathbb{T}^{\dagger}\geq\delta^{\eta/2}(\#\mathbb{T})$. In particular, by Remark 2.11 this means that $\mathbb{T}^{\dagger}$ and each set $\mathbb{T}^{\dagger}_{s_{i}},\ i=1,\ldots,N$ satisfies the Convex Wolff Axioms with error $\delta^{-2\eta}$.

### v2

@@ 4.1 Arrangements of rectangular prisms
i.e. $\bigcup_{\mathcal{R}}Y(R)$ has discretized Assouad dimension $\geq 3-\varepsilon$ at scale $\delta$ and scale separation $\delta^{-\tau_{i}}$. If this occurs, then we halt the iteration and conclude the proof.
Case (B). Next, suppose that Conclusion (B) from Lemma 4.2 holds. Define $s_{i+1}=s_{i}^{\prime}$ and $t_{i+1}=t_{i}^{\prime}$ (here $s_{i}^{\prime}$ and $t_{i}^{\prime}$ are the output of Lemma 4.2, Conclusion (B)). Let $\mathcal{R}_{i+1}$ be the set of $s_{i+1}\times t_{i+1}\times 1$ prisms, with shading $Y_{i+1}=Y_{i}^{\prime}$, as described by Lemma 4.2, Conclusion (B). We have that $\mathcal{R}_{i+1}$ satisfies the Convex Wolff Axioms with error $\delta^{-\tau_{i+1}}$, and $Y_{i+1}$ is $\delta^{\tau_{i+1}}$ dense. Furthermore, the shading $Y_{i+1}$ satisfies (4.27) with $i+1$ in place of $i$. In particular, $\mathcal{R}_{i+1}$ and $Y_{i+1}$ satisfy our induction hypotheses (b), (c), and (d) above.
Case (B.i). Suppose that Conclusion (B) holds and that $t_{i+1}=1$. Recall that $s_{i+1}\leq\delta^{\varepsilon/2}.$ Thus we can apply Lemma 4.4 with $s_{i+1}$ in place of $\delta$. We have
@@
We can select $\tau_{i+1}$ and $\delta_{0}$ sufficiently small so that the LHS of (4.28) has size at least $s_{i+1}^{\varepsilon}|B(0,1)|$. Since the set on the LHS of (4.28) is contained in $B(0,1)$, we conclude that $\bigcup_{\mathcal{R}}Y(R)$ has discretized Assouad dimension $\geq 3-\varepsilon$ at scale $\delta$ and scale separation $\delta^{-\varepsilon/2}\geq\delta^{-\eta}.$ If this occurs, then we halt the iteration and conclude the proof.
Case (B.ii). Suppose $t_{i+1}<1$; then $s_{i+1}/t_{i+1}=s_{i}/t_{i}$, so in particular $s_{i+1}\leq\delta^{\varepsilon}t_{i+1}$. This means that the induction hypothesis (a) also holds, and thus we have verified all of the properties required to continue the iteration, with $i+1$ in place of $i$. Note that $s_{i+1}\geq\delta^{-\varepsilon^{2}/4}s_{i}$; $t_{i+1}\geq\min(1,\delta^{\varepsilon}s_{i+1})$, and the iterative process halts if $t_{i+1}=1$. We conclude that this iterative process will halt after at most $N=\lfloor 4/\varepsilon^{2}\rfloor$ steps. ∎
@@ 4.2 Arrangements of tubes that cluster into rectangular prisms
In Section 4.1, we showed that unions of planks have discretized Assouad dimension close to 3. In this section we will use these results to show that if an arrangement of $\delta$-tubes clusters into planks, then the union of these tubes must have discretized Assouad dimension close to 3. This is made precise in Lemmas 4.7 and 4.9 below.
As a first step, the following lemma says that if $\mathbb{T}$ is a set of $\delta$ tubes of cardinality $>\!\!>\rho/\delta$, all of which are contained in a common $\delta\times\rho\times 1$ prism, then these tubes must cluster into $\delta\times\omega\times 1$ planks (for some $\delta<\!\!<\omega\leq\rho$), and the union of these tubes must fill out a substantial fraction of each $\delta\times\omega\times 1$ plank. Later, we will show that the existence of such (mostly full) $\delta\times\omega\times 1$ planks will force $\bigcup_{\mathbb{T}}T$ to have discretized Assouad dimension close to 3.
@@
For all $\varepsilon>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $\rho\in[20\delta,1]$ and let $K\geq 1$. Let $(\mathbb{T},Y)_{\delta}$ be a set of essentially distinct $\delta$-tubes, with #$\#\mathbb{T}\geq K\rho/\delta$, and $Y$ a uniformly $\delta^{\eta}$-dense shading. Let $R_{0}$ be a $10\delta\times\rho\times 1$ rectangle, and suppose each $T\in\mathbb{T}$ is contained in $R_{0}$.
@@
where the maximum is taken over all pairs of tubes $T,T^{\prime}$ intersecting $x$. We define $\operatorname{L-spread}(x)=0$ if at most one tube intersects $x$. Let $C$ be a large number to be determined below. We claim that there exists a choice of $\theta\in[\frac{K\delta}{C},\rho]$ so that the shading
@@
(Note that the shading $Z(T)$ is not necessarily a subset of $Y(T)$). We justify this claim as follows. Since there are at most $10\log(C/\delta)$ dyadic values of $\theta$ in the interval $[\frac{K\delta}{C},\rho]$, if we could not select a $\theta\in[\frac{K\delta}{C},\rho]$ so that (4.31) holds, then (provided $C$ is selected sufficiently large) the shading
@@
satisfies $\sum_{\mathbb{T}}|Z^{\prime}(T)|\geq\frac{1}{100}\delta^{2}(\#\mathbb{T})$. But for each $x\in\mathbb{R}^{3}$, the set of tubes $T\in\mathbb{T}$ with $x\in Z(T)$ are contained in the $10\delta$-neighborhood of a plane (since they are contained in $R_{0}$), and also make angle $\leq\frac{K\delta}{C}$ with a common unit vector; since the tubes in $\mathbb{T}$ are essentially distinct, we conclude that $O(K/C)$ tubes $T\in\mathbb{T}$ can satisfy $x\in Z^{\prime}(T)$, and hence
@@
$\mathbb{T}_{1}=\big{\{}T\in\mathbb{T}\colon|Z(T)|\geq(C|\log\delta|)^{-1}|T|\big{\}}.$ (4.33)
Then (4.31) guarantees that #$\#\mathbb{T}_{1}\gtrapprox\#\mathbb{T}$. For each $T_{1}\in\mathbb{T}_{1}$, let $x_{2},\ldots,x_{N},\ N\gtrapprox\theta/\delta$ be a set of $\delta/\theta$-separated points from $Y_{1}(T_{1})$. For each $x_{i}$, select $T_{i}\in\mathbb{T}$ with $x_{i}\in T_{i}$ and $\angle(T_{1},T_{i})\in[\theta,2\theta)$. An $L^{2}$ argument (see e.g. (4.15)) plus the fact that $|Y(T_{i})|\geq\delta^{\eta}|T|$ shows that
@@
On the other hand, the set on the LHS of (4.34) is contained in $R_{0}\cap N_{2\theta}(T_{1})$, which in turn is contained in a rectangular prism of dimensions $10\delta\times 4\theta\times 1$; denote this prism by $R(T_{1})$. We have just shown that
@@
Finally, let $\omega=4\theta$ and let $\mathcal{R}_{1}=\{R(T_{1})\colon T_{1}\in\mathbb{T}_{1}\}$. We claim that (4.29) holds for $\mathcal{R}_{1}$. Indeed, this follows from the fact that $T_{1}\subset R(T_{1})$ for each $T_{1}\in\mathbb{T}_{1}$, and #$\#\mathbb{T}_{1}\gtrapprox\#\mathbb{T}$. By (4.35) we see that (4.30) holds for each $R\in\mathcal{R}_{1}$, provided we select $\eta$ and $\delta_{0}$ sufficiently small. To conclude the proof, greedily select a set $\mathcal{R}\subset\mathcal{R}_{1}$ of essentially distinct rectangular prisms; we can perform this selection so that (4.29) continues to hold. ∎
The following lemmas says that if $\mathbb{T}$ clusters into planks, then it must have discretized Assouad dimension close to 3.
@@
For all $\varepsilon>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $\mathbb{T}$ be a set of essentially distinct $\delta$-tubes that satisfies the Convex Wolff Axioms with error $\delta^{-\eta}$. Let $Y$ be a uniformly $\delta^{\eta}$-dense shading. Then at least one of the following must occur:
@@
    $\bigcup_{\mathbb{T}}Y(T)$ has discretized Assouad dimension at least $3-\varepsilon$ at scale $\delta$ and scale separation $\delta^{-\eta}$.
@@
    There is a set $\mathbb{T}^{\prime}\subset\mathbb{T}$ with #$\#\mathbb{T}^{\prime}\geq\frac{1}{2}(\#\mathbb{T})$ so that for every $\delta\leq\rho\leq 1$ and every $\delta\times\rho\times 1$ rectangle $R$, at most $\delta^{-\varepsilon}(\rho/\delta)$ tubes from $\mathbb{T}^{\prime}$ are contained in $100R$.
@@
Since $\mathbb{T}$ satisfies the Convex Wolff Axioms, we have that at most $100^{3}\delta^{-\eta}(\rho\delta)(\#\mathbb{T})$ tubes are contained in each $100\delta\times 100\rho\times 100$ rectangle. If #$\#\mathbb{T}$ is substantially larger than $\delta^{-2}$, then the estimate from Conclusion (B) is stronger.
@@
We say a $100\delta\times 100\rho\times 100$ rectangle is “heavy” if it contains more than $\delta^{-\varepsilon}(\rho/\delta)$ tubes from $\mathbb{T}$. If fewer than $\frac{1}{2}(\#\mathbb{T})$ of the tubes are contained in heavy rectangles, let $\mathbb{T}^{\prime}$ be the set of tubes that are not contained in heavy rectangles. Then Conclusion (B) holds, and we are done.
Suppose instead that at least $\frac{1}{2}(\#\mathbb{T})$ tubes are contained in heavy rectangles. Since the tubes in $\mathbb{T}$ are essentially distinct, each $100\delta\times 100\rho\times 100$ rectangle can contain at most $O\big{(}(\rho/\delta)^{2}\big{)}$ tubes from $\mathbb{T}$. Thus if a $100\delta\times 100\rho\times 100$ rectangle is heavy, then we must have $\rho\gtrsim\delta^{1-\varepsilon}$. Let $\mathcal{R}_{0}$ denote this set of heavy rectangles (note that the rectangles in $\mathcal{R}_{0}$ can have different dimensions). Let $\varepsilon_{1}$ be a small quantity to be chosen below. Apply Lemma 4.6 to each rectangle $R_{0}\in\mathcal{R}_{0}$, with $\varepsilon_{1}$ in place of $\varepsilon$, and $K=\delta^{-\varepsilon}$. We obtain a number $\omega(R_{0})\gtrsim\delta^{1-\varepsilon}$ and a set $\mathcal{R}(R_{0})$ of $\delta\times\omega(R_{0})\times O(1)$ rectangles. We have
@@
Since $\mathbb{T}$ satisfies the Convex Wolff Axioms with error $\delta^{-\eta}$, by (4.36) we have that the set (i.e. not multi-set) $\mathbb{T}_{2}$ satisfies the Convex Wolff Axioms with error $\lessapprox\delta^{-\eta}$. Finally, for each $T\in\mathbb{T}_{2}$, let $R(T)$ be some rectangle in $\mathcal{R}_{2}$ with $T\subset R$. Define the multi-set =$\mathcal{R}=\{R(T)\colon T\in\mathbb{T}_{2}\}$, i.e. $\mathcal{R}$ is a multi-set with the same cardinality as $\mathbb{T}_{2}$. By Remark 2.7, we conclude that $\mathcal{R}$ satisfies the Convex Wolff Axioms with error $\lessapprox\delta^{-\eta}$.
@@
By Lemma 4.6, we have $|Y(R)|\geq\delta^{\varepsilon_{1}}|R|$ for each $R\in\mathcal{R}$. To conclude the proof, apply Proposition 4.5 with $\varepsilon$ as above, $s=\delta$, and $t=\omega_{0}$; to ensure that Proposition 4.5 can be applied, we must select $\varepsilon_{1}$ sufficiently small depending on $\varepsilon$ (our choice of $\varepsilon_{1}$ depends on the quantity “$\eta$” from Proposition 4.5), and then $\eta$ sufficiently small depending on $\varepsilon_{1}$. Since $\bigcup Y(T)\subset\bigcup Y(R),$ we conclude that Conclusion (A) holds. ∎
Lemma 4.7 says that either a union of $\delta$-tubes $\mathbb{T}$ has large discretized Assouad dimension, or the tubes in $\mathbb{T}$ cannot concentrate into $\delta\times\rho\times 1$ rectangular prisms. The next lemma is similar, except the latter conclusion is strengthened: either a union of $\delta$-tubes has large discretized Assouad dimension, or these tubes cannot concentrate into $s\times t\times 1$ rectangular prisms, for any $\delta\leq s\leq t\leq 1$. To prove this result, we apply Lemma 4.7 to the $s$-thickening of the tubes in our arrangement, for many different values of $s$. The precise statement is as follows.
@@
For all $\varepsilon=1/N>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $(\mathbb{T},Y)_{\delta}$ be a set of essentially distinct $\delta$-tubes that satisfies the Convex Wolff Axioms with error $\delta^{-\eta}$, and let $Y$ be a uniformly $\delta^{\eta}$-dense shading. Then at least one of the following must occur:
@@
    $\bigcup_{\mathbb{T}}Y(T)$ has discretized Assouad dimension at least $3-\varepsilon$ at scale $\delta$ and scale separation $\delta^{-\eta}$.
@@
    There is a set $\mathbb{T}^{\prime}\subset\mathbb{T}$ with #$\#\mathbb{T}^{\prime}\geq\delta^{\varepsilon}(\#\mathbb{T})$ so that the following holds. For every number $\delta\leq s\leq 1$ of the form $\delta^{i\varepsilon}$, there is a uniform partitioning cover $\mathbb{T}_{s}$ of $\mathbb{T}^{\prime}$ (see Definition 2.10), where $\mathbb{T}_{s}$ is a set of essentially distinct $s$-tubes. Furthermore, for every $t\in[s,1]$ and every $s\times t\times 1$ rectangular prism $R$, at most $\delta^{-\varepsilon}\frac{t}{s}\frac{(\#\mathbb{T}^{\prime})}{(\#\mathbb{T}_{s})}$ tubes from $\mathbb{T}^{\prime}$ are contained in $100R$.
@@
Our first step is to regularize the set $\mathbb{T}$ at scales of the form $s_{i}=\delta^{i/N}$, with $i=1,\ldots,N$ (we do not need to consider $s_{0}=1$, since in that case there is nothing to prove). After repeated pigeonholing, we can select a set $\mathbb{T}^{\dagger}\subset\mathbb{T}$ with #$\#\mathbb{T}^{\dagger}\geq(100\log(1/\delta))^{-N}(\#\mathbb{T})$ and sets of $s_{i}$-tubes $\mathbb{T}_{s_{i}},\ i=0,\ldots,N$, so that $\mathbb{T}_{s_{N}}=\mathbb{T}^{\dagger}$, and for each $i=1,\ldots,N-1$, the set $\mathbb{T}_{s_{i}}$ is a uniform partitioning cover of $\mathbb{T}_{s_{i+1}}$ (and hence also a uniform partitioning cover of $\mathbb{T}_{s_{j}}$ for all $j>i$). See e.g. Lemma 3.4 from [11] for an explicit description of this iterated pigeonholing process. We will choose $\delta_{0}$ sufficiently small so that #$\#\mathbb{T}^{\dagger}\geq\delta^{\eta/2}(\#\mathbb{T})$. In particular, by Remark 2.11 this means that $\mathbb{T}^{\dagger}$ and each set $\mathbb{T}_{s_{i}},\ i=1,\ldots,N$ satisfies the Convex Wolff Axioms with error $\delta^{-2\eta}\leq s_{i}^{-2\eta N}$. For each index $i$ and each $T_{s_{i}}\in\mathbb{T}_{s_{i}}$, define the shading $Y_{s_{i}}(T_{s_{i}})=T_{s_{i}}\cap N_{s_{i}}(Y(T))$, where $T$ is a tube from $\mathbb{T}[T_{s_{i}}]$. In particular, $|Y_{s_{i}}(T_{s_{i}})|\gtrsim\delta^{\eta}|T_{s_{i}}|\gtrsim s_{i}^{\eta N}|T_{s_{i}}|$, i.e. $Y_{s_{i}}$ is a uniformly $s_{i}^{\eta N}$-dense shading of $\mathbb{T}_{s_{i}}$.